OpenStudy (anonymous):

Write this in the Standard Vertex form: y=x^2-6x+14

OpenStudy (anonymous):

Complete square, can you?

OpenStudy (anonymous):

oh ok thanks yeah, i can complete squares :)

OpenStudy (anonymous):

Would you like me to double check your work?

OpenStudy (anonymous):

ok, i'll go do it now then

OpenStudy (anonymous):

like, redo it haha

OpenStudy (anonymous):

y=(x-3)^2=5

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

y= (x-3)² + 5 So you know (h, k) !